### v2

# ReMatch: Retrieval Enhanced Schema Matching with LLMs

Abstract.
Schema matching is a crucial task in data integration, involving the alignment of a source schema with a target schema to establish correspondence between their elements. This task is challenging due to textual and semantic heterogeneity, as well as differences in schema sizes. Although machine-learning-based solutions have been explored in numerous studies, they often suffer from low accuracy, require manual mapping of the schemas for model training, or need access to source schema data which might be unavailable due to privacy concerns. In this paper we present a novel method, named ReMatch, for matching schemas using retrieval-enhanced Large Language Models (LLMs). Our method avoids the need for predefined mapping, any model training, or access to data in the source database. Our experimental results on large real-world schemas demonstrate that ReMatch is an effective matcher. By eliminating the requirement for training data, ReMatch becomes a viable solution for real-world scenarios.
Schema Matching, Large Language Models (LLMs), RAG, Machine Learning, Ranking, Data Integration
ccs: Information systems Mediators and data integrationccs: Computing methodologies Natural language processing
1. Introduction
Figure 1. Overview of the ReMatch method.
Schema matching is a fundamental task in data management and integration, involving the identification of semantic correspondences between elements of two or more database schemas (Madhavan et al., 2001), regardless of differences in naming, structure, or data type. The matching task is challenging since schemas are designed using different perspectives and terminologies. This semantic heterogeneity can lead to ambiguous mappings where schema elements have the same name but different meanings, or different names but the same meaning.
Human schema matching, a manual and time-consuming process, requires significant effort from skilled individuals. This can be expensive and impractical, particularly in large-scale projects. Furthermore, human matchers are prone to errors and inconsistencies due to cognitive biases and fatigue (Shraga et al., 2021). As a result, the automation of schema matching has become a major focus within the AI and database-oriented research community over the years.
In recent years, LLMs have achieved significant advancements across many challenging tasks that require a deep understanding of semantics. These models have shown an impressive ability to generalize to new tasks without any task-specific fine-tuning, even in areas significantly divergent from the ones they were originally trained on, including various data related tasks (Narayan et al., 2022; Zhang et al., 2023a).
In this paper, we present a new approach that unlike previously proposed machine learning (ML) methods, circumvents the need for predefined mapping, model training, or access to data in the source database. Our approach, which we call ReMatch, improves the task of schema matching by using retrieval-enhanced LLMs. Fig. 1 illustrates our proposed approach.
Our contributions in this study are as follows: (1) Introducing a new method for the task of schema matching, which allows for scalable and accurate matching results, without any model training or access to labeled data. (2) Proposing a mechanism to reduce the search space of the target schema for efficient candidate generation. (3) Exploiting the generative abilities of LLMs to perform semantic ranking between two schemas, in alignment with human matchers. (4) Providing a complete mapping between two important healthcare schemas, aiming to address the lack of real-world evaluation datasets in the field. To the best of our knowledge, this is currently the largest publicly available schema matching dataset.
2. Background and Related Work
Large Language Models (LLMs)  Generative language models are trained to generate human-like text and can be fine-tuned for a variety of tasks, but also can be reused for a variety of tasks, with no additional training (Radford et al., 2018; Brown et al., 2020; Liu et al., 2023; Mirchandani et al., 2023). Embedding models like BERT or Ada (Neelakantan et al., 2022) provide contextual embeddings that have significantly improved performance on a wide range of NLP tasks (Sun et al., 2019; Minaee et al., 2021). Moreover, the representation of text via embeddings allows for efficient and accurate passage retrieval, using semantic similarity (Neelakantan et al., 2022; Lewis et al., 2020).
LLMs have been applied to various data preprocessing tasks, including error detection, data imputation, entity matching, as well as schema matching (Mirchandani et al., 2023; Narayan et al., 2022). These models have shown promising results, however, these approaches still faced challenges related to computational expense and inefficiency, as well as lacking evaluation on data that represents real-world schema matching challenges.
Schema Matching  Schema matching involves identifying correspondences between different data schemas. Traditionally, schema matching was done manually, however, this approach strongly relies on the expertise of human matchers (Dragisic et al., 2016). This has led to development of automated schema matching, primarily using ML algorithms (Do and Rahm, 2002; Bernstein et al., 2011; Madhavan et al., 2001). Much of the work on ML-based schema matching was limited to basic schemas, with most works dealing only with toy datasets (Do and Rahm, 2002; Gal et al., 2021; Mudgal et al., 2018; Shraga et al., 2020; Li et al., 2020), or focused primarily on entity matching only. This stemmed from a combination of lack of high-quality datasets (Koutras et al., 2020), as well as relatively poor results of past methods.
Recent work, such as SMAT (Zhang et al., 2021) and LSM (Zhang et al., 2023b), leverages advancements in natural language processing to achieve semantic mappings between source and target schemas. While these methods improve previous results, they still require extensive data tagging, limiting their practicality in real-world applications.
Finally, the majority of previous methods dealt with schema-matching as a binary classification task over $\mathcal{A}_{1}\times\mathcal{A}_{2}$. While simplifying training and evaluation, in normalized schemas positive labels scale by $O(n)$, whereas negative labels scale by $O(n^{2})$, which is problematic when inference is expensive, as in the case of LLMs.
3. ReMatch
Problem Statement  Given a source schema $\mathcal{S}_{1}$ and a target schema $\mathcal{S}_{2}$, with sets of tables $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ and sets of attributes $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively, the schema matching task involves finding a mapping between $(\mathcal{A}_{1},\mathcal{T}_{1})\in\mathcal{S}_{1}$ to $(\mathcal{A}_{2},\mathcal{T}_{2})\in\mathcal{S}_{2}$. The matches between their attributes are captured by a relation match: $\mathcal{P}(\mathcal{A}_{1})\times\mathcal{P}(\mathcal{A}_{2})$. An element $(A_{1},A_{2})\in$ match defines a matching pair possibly representing the same information in the schemas. If $\lvert A_{1}\rvert=1$ and $\lvert A_{2}\rvert=1$, we call the match an elementary match or 1:1 match. Otherwise, we refer to it as a complex match or m:n match.
Formally, we want to find a function $\Psi:\mathcal{A}_{1}\rightarrow\mathcal{P}(\mathcal{A}_{2})$, such that $\forall a\in\mathcal{A}_{1},\>\forall a^{\prime}\in\Psi(a),\>(a,a^{\prime})\in match$. In other words, $\Psi(a)=A^{\prime}\subseteq\mathcal{A}_{2}$, should contain all relevant matches for $a$.
A possible simplification of this problem can be made by limiting $\Psi$ to be a $N\times K,\>N=|\mathcal{A}_{1}|,K\in\mathbb{N}$, matrix of elements from $\mathcal{A}_{2}$, denoted as $\Psi_{K}$. In the case of elementary matches, or m:1 complex matches, the goal from above now becomes to maximize the $accuracy@K$ metric, i.e., to maximize:
(1) $\frac{1}{N}\sum^{N}\mathbf{1}_{\{\exists a^{\prime},\>(a,a^{\prime})\in match,\>a^{\prime}\in\Psi_{K}(a)\}}$
Method Description   Algorithm 1 shows the pseudocode for our method. Given a source schema $\mathcal{S}_{1}$ with a set of tables $\mathcal{T}_{1}$ and a set of attributes $\mathcal{A}_{1}$, and a target schema $\mathcal{S}_{2}$ with a set of tables $\mathcal{T}_{2}$ and a set of attributes $\mathcal{A}_{2}$:
(1) Target schema tables and source schema attributes are transformed into two corpora of structured documents, $\mathcal{C}_{t}$ and $\mathcal{C}_{s}$, respectively. Each target schema table and source schema attribute is represented as a structured document consisting of descriptive paragraphs. The title of the document is the table’s name, and the opening paragraph provides an overview of the table’s purpose and characteristics. The subsequent paragraphs detail the set of attributes serving as the table’s primary key, the set of attributes referring to other tables (foreign keys), and the rest of the attributes belonging to this table, respectively. Each attribute is followed by its data type and a textual description. In the case of attribute documents, the specific attribute is highlighted above the title.
(2) For each attribute in the source schema, we search for the top $J$ documents that represent candidate tables from the target schema. This step ensures that only the most relevant documents are considered for matching. To facilitate the retrieval of candidate tables, we utilize a text embedding model to encode each source attribute’s document, and the corpus of target table documents. These embeddings serve as a basis for measuring semantic similarity, enabling the retrieval of candidate tables. Finally, for every source table $t_{i}\in\mathcal{T}_{1}$, we create a set of all top $J$ candidate tables $\mathcal{T}_{c}$ retrieved.
(3) In the last step, the LLM is tasked with selecting the top $K$ most similar target attributes from the set of retrieved tables $\mathcal{T}_{c}$ identified earlier. The model assesses the context provided by the document representations, and yields a ranked list of $K$ potential matches for each attribute in the source schema.
Inputs :  a source schema $\mathcal{S}_{1}$, $\mathcal{T}_{1}$, $\mathcal{A}_{1}$, with all its textual descriptions; a target schema $\mathcal{S}_{2}$,$\mathcal{T}_{2}$, $\mathcal{A}_{2}$ with all its textual descriptions; $K$,$J\in\mathbb{N}$; an embedding model $\Phi$; a generative LLM $\mathcal{F}$.
Outputs : a $N\times K$ matrix of attribute candidates from $\mathcal{A}_{2}$, denoted as $\Psi_{K}$.
1 $\mathcal{C}_{s}\leftarrow\emptyset$, $\mathcal{C}_{t}\leftarrow\emptyset$, $\Psi_{K}\leftarrow\{\}$
/* lines 2-4 are performed once, as a pre-processing step. */
2 $\mathcal{C}_{t}\leftarrow TableToDoc(\mathcal{S}_{2},\mathcal{T}_{2},\mathcal{A}_{2})$
3 $\mathcal{C}_{s}\leftarrow AttributeToDoc(\mathcal{S}_{1},\mathcal{T}_{1},\mathcal{A}_{1})$
4 $\tilde{\mathcal{C}_{t}}\leftarrow\Phi(\mathcal{C}_{t})$
5 foreach $t_{i}\in\mathcal{T}_{1}$ do
6       $\mathcal{T}_{c}\leftarrow\{\}$
7       foreach $a_{k}\in\mathcal{A}_{1}[t_{i}]$ do
8             $\tilde{a}_{emb}\leftarrow\Phi(\mathcal{C}_{s}[t_{i};a_{k}])$
9             $scores\leftarrow Similarity(\tilde{a}_{emb},\tilde{\mathcal{C}_{t}})$
10             $top_{j}\leftarrow\mathrm{argmaxes}(\{scores_{1},\ldots,scores_{n}\})[1:j]$
11             $\mathcal{T}_{c}\leftarrow\mathcal{T}_{c}\cup\mathcal{C}_{t}[top_{j}]$
12       end foreach
13      $\Psi_{K}\leftarrow CreateTopkMapping(t_{i},\mathcal{C}_{s}[t_{i}],\mathcal{T}_{c},\mathcal{C}_{t}[\mathcal{T}_{c}],\mathcal{F})$
14 end foreach
Algorithm 1 ReMatch
4. Evaluation
4.1. Dataset Creation
To evaluate our method we used two primary datasets, both involving mappings between healthcare database schemas.
MIMIC-III to OMOP  For the first dataset, we created a mapping between the schema of MIMIC-III (Johnson et al., 2016), a public database containing deidentified records from the Beth Israel Deaconess Medical Center, and The Observational Medical Outcomes Partnership Common Data Model (OMOP) [^1] https://www.ohdsi.org/data-standardization/, an open-source data standard. The mapping was created manually by a domain expert, aided by the mapping created in (Paris et al., 2021). If no matching attribute in OMOP could be found, the attribute was assigned NA. For convenience we will refer to this dataset as MIMIC.
OMAP Benchmark Synthea Dataset  Following (Narayan et al., 2022) we used the Synthea (Walonoski et al., 2018) to OMOP dataset from the OMAP benchmark (Zhang et al., 2021). The dataset contains a partial mapping of the schema for Synthea (the source), a synthetic healthcare dataset, to a partial subset of relevant OMOP attributes (the target). For convenience we will refer to this dataset as Synthea.
It is important to note that although both datasets are similar, the first maps the entire MIMIC-III schema to OMOP, while Synthea contains only partial mappings. To the best of our knowledge, the MIMIC dataset is the largest single-schema to single-schema dataset published. Table 1 contains statistics detailing both datasets. The full MIMIC dataset is publicly available at https://github.com/meniData1/MIMIC_2_OMOP.
4.2. Experiments
We established the performance of ReMatch and SMAT (Zhang et al., 2021), the previous state-of-the-art (SOTA) method, on MIMIC and Synthea.
As our method treats schema matching as a retrieval problem, the most appropriate metric is accuracy@$K$, as defined in Eq. 1. Unlike the F1-score, standard accuracy@$K$ is not well defined for m:n matches. To deal with this, we limited ourselves to evaluating 1:1 or m:1 relations.  [^2] It is worth noting that for a dataset with only 1:1 and m:1 mappings, accuracy@1 is equivalent to F1-score when using argmax, instead of a threshold, for prediction (since it trivially reflects both the precision and the recall simultaneously).
There are multiple dimensions we would like to optimize. The first is maximizing accuracy@$K$. We also seek lower values of $K$, as our aim is to assist human schema matching, and minimize cognitive overload (Bernstein et al., 2011; Chernev et al., 2015). Finally, we would like to minimize the number of target tables, reducing prompt size and costs.
Evaluating ReMatch  For all experiments we used GPT-4 (OpenAI, 2023), specifically GPT-4-1106, and Ada2 (Greene et al., 2022; Neelakantan et al., 2022) as our embedding model.
We performed a grid search over various values of $J$ and $K$, also allowing us to select hyperparameters for further evaluations. As an ablation study, we evaluated the performance on MIMIC when only the names of the tables and attributes are used for retrieval and generation. We also evaluated a version with no retrieval stage, i.e., for each table in the source schema we used the entire target schema. Finally, we ran ReMatch multiple times on each of the setups in the ablation study, to verify stability across generations.
Evaluating ReMatch with Guidance  A human matcher may determine that the suggestions are not good enough, or decide to provide information they already have. In our evaluation we investigated a mechanism, inspired by the relevance feedback technique (Rocchio, 1971). To do so, we selected a mapping of the form $(T_{1},a_{1})\rightarrow(T_{2},a_{2})$ for a single attribute from each table. Specifically, we used the SUBJECT_ID attribute where available (19/25 tables), and an attribute containing unique identifiers otherwise. We then automatically included $T_{2}$ in the set of retrieved tables, $\mathcal{T}_{c}$, and provided the LLM with this mapping in the prompt.
Evaluating SMAT  Our method’s performance was evaluated against SMAT (Zhang et al., 2021), a previously SOTA non-LLM model. For SMAT’s training we used the default hyperparameters: {Learning Rate: 0.8, Batch Size: 64, Epochs: 30}, with early stopping, and 10% of the training data was used as a validation set to avoid over-fitting. Models were trained on a single Nvidia A-100 GPU.
We tried two setups, 80% training and 20% testing, and vice versa. The splits were stratified across table names. We also extended the evaluation to include accuracy@$K$. We calculated the accuracy@$K$ by determining whether the correct match received one of the $K$ highest classification scores. Binary predictions were given by using the optimal threshold for the validation set.
Table 1. Dataset statistics for MIMIC and Synthea. #Columns and #Tables refer to the number of columns and tables used in the dataset. #Mapped Columns refers to the number of unique columns mapped. #Null Mappings refers to source columns without a mapping (mapped to NA).
Dataset #Columns #Tables #Mapped Columns #Null Mappings
MIMIC 268 25 156 112
OMOP (MIMIC) 425 38 95 -
Synthea 38 8 105 -
OMOP (Synthea) 67 8 67 -
Table 2. Grid search results on MIMIC.
Retrieved Documents Acc@1 Acc@2 Acc@3 Acc@5 Acc@7 Avg #T
$J=1$ 0.424 0.589 0.644 0.709 0.697 2.44
$J=2$ 0.425 0.541 0.638 0.729 0.758 4.68
$J=3$ 0.321 0.477 0.533 0.657 0.733 6.64
$J=5$ 0.336 0.425 0.504 0.657 0.754 9.88
$J=7$ 0.317 0.399 0.414 0.616 0.71 12.84
Table 3. Grid search results on Synthea. $J=\infty$ means skipping the retrieval step.
Retrieved Documents Acc@1 Acc@3 Acc@5 Avg #T
$J=1$ 0.562 0.6 0.6 1.125
$J=2$ 0.505 0.581 0.743 2.625
$J=\infty$ 0.438 - 0.924 8
Table 4. Results on MIMIC for ablation study with standard deviations across runs. Baseline is the full ReMatch method, names-only refers to skipping the document creation step, and $J=\infty$ means skipping the retrieval step.
Variation Acc@1 Acc@5 Avg #T
$J=1$ baseline 0.424 $\pm$ 0.017 - 2.44
$J=2$ baseline - 0.729 $\pm$ 0.023 4.68
$J=1$ names-only 0.396 $\pm$ 0.005 - 1.48
$J=2$ names-only - 0.503 $\pm$ 0.008 3.28
$J=\infty$ 0.311 $\pm$ 0.0157 0.518 $\pm$ 0.03 38
Table 5. Results for MIMIC with guidance provided. Relative improvement over baseline (in percentage) in bold.
Variation Acc@1 Acc@5 Avg #T
$J=1$ 0.539/20.38% 0.783/10.41% 2.84
$J=2$ 0.459/7.89% 0.765/4.86% 4.92
4.3. Results
ReMatch Results The results of the grid search for MIMIC can be found in Table 2. We found $\{(J=1,\>K=1),\>(J=2,\>K=5)\}$ to be the best balance points between the three optimization dimensions for the MIMIC dataset. Interestingly, accuracy@$K$ seems to perform best for $J\leq 2$. This may imply that the retrieval is efficient, and additional tables only add ‘noise’ to the matching prompt.
The results for the grid search for Synthea are presented inTable 3. While $\{(J=1,\>K=1),\>(J=2,\>K=5)\}$ still appear to work well, for Synthea we found that not using retrieval yielded optimal results for accuracy@5. This may be explained by the small size of the dataset, with only 8 tables in the target.
The results of the ablation study are shown in Table 4. We found that skipping the retrieval step yields inferior results, in alignment with the results from the grid search. We also found that using only the names of the tables and attributes, with no descriptions, results in a significant decrease. Additionally, Table 4 demonstrates that the method’s outputs are stable.
Adding guidance to ReMatch indeed improves performance, as can be seen in Table 5. The largest improvements are for $J=1$, achieving the highest overall scores for MIMIC. This is consistent with the rest of the results, with more efficient retrieval being correlated with better matching accuracy.
SMAT Results  SMAT’s performance was validated using the F1-score, with an F1-score of 0.48276 for the 80-20 train-test split on the Synthea dataset, indicating successful model training when compared to the original values reported by Zhang et al.
When using 80% of the data for training, SMAT achieved a perfect accuracy@5 score (accuracy@5 = 100%) on the remaining 20% of the attributes. Unfortunately, in real-world scenarios only a small portion of mappings can be available, since mapping 80% manually removes the need for automatic mapping. In the much more realistic split, with only 20% of the data, SMAT achieved significantly lower results than ReMatch on both datasets, as is emphasized in Fig. 2. Unlike ReMatch, the results were significantly higher on Synthea, when compared to MIMIC dataset, for both splits. This finding is aligned with our expectations about the difficulty of MIMIC, and demonstrates the robustness of ReMatch.
ACC@1ACC@5000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.80.0770.0770.0770.0770.3610.3610.3610.3610.4240.4240.4240.4240.7290.7290.7290.729SMATReMatch
(a) Performance on MIMIC.
ACC@1ACC@5000.50.50.50.511110.2710.2710.2710.2710.6940.6940.6940.6940.5620.5620.5620.5620.9240.9240.9240.924SMATReMatch
(b) Performance on Synthea.
Figure 2. Comparison of the different models’ performance. SMAT was trained and evaluated on 20%, 80% of the data, after removing all null mappings. ReMatch was evaluated on the entire dataset, with no guidance, and with nulls. Optimal setup from grid search is shown for ReMatch.
Table 6. Results for SMAT on both datasets.
Variation Acc@1 Acc@5 F1
SMAT 80-20 0.454 1.0 0.102
SMAT 20-80 0.077 0.361 0.084
Variation Acc@1 Acc@5 F1
SMAT 80-20 0.875 1.0 0.483
SMAT 20-80 0.271 0.694 0.241
5. Conclusions
In this work we introduced ReMatch, a scalable and effective method for matching schemas using retrieval-enhanced LLMs, designed to complement and aid human matchers throughout their work. ReMatch avoids the need for predefined mapping, any model training, or access to data in the source database. Instead, it exploits the generative abilities of LLMs to perform semantic ranking between two schemas, in alignment with a human matching process. We also provided a new large dataset that will hopefully aid further research.
In future work, we plan to include the explicit use of type constraints, foreign keys, and primary keys, as well as enhanced guidance mechanisms and enrichment of table and column descriptions using the source schema data, where accessible without privacy or security constraints.